Consider the following projects X and Y where the firm can choose only once. Project X costs $600
and has cash flows of $400 in each of the next two years. Project B also costs $ 600 and generates
cash flows $500 and $275 for the next two years, respectively. Sketch a net present value profile
(graphs) for each of these projects. For graphs you may use approximation

Which project should the firm choose if the cost of capital is 10 percent? What if the cost of capital
is 25 percent?
Answer #1
A. Cost of capital = 10%
NPV of Proj X = NPV(Rate, CF1,CF2) + CF0 = NPV(10%,400,400) -600 = $94
NPV of Proj Y = NPV(10%,500,275) - 600 = $82

Based on NPV, Proj X is better due to higher NPV


B. Cost of capital = 25%
NPV of Proj X = NPV(Rate, CF1,CF2) + CF0 = NPV(25%,400,400) -600 = ($24)
NPV of Proj Y = NPV(25%,500,275) - 600 =($24)

Based on NPV, Both projects should be rejected.